Find an expression for a square matrix A satisfying A 2 = In, where In is the n × n identity matrix


1
Expert's answer
2021-05-18T14:13:19-0400

For any matrix A = I - 2 * v*vT, where v is a unit vector of dimension n, and I is identity matrix

A2 = A * A = (I - 2 v vT) * (I - 2 v vT) =

I - I * (2 v vT) - (2 v vT) * I + (2 v vT) * (2 v vT) =

I - 4 v vT + 4 v vT v vT =

I - 4 v vT + 4 v (vT v) vT =

I - 4 v vT + 4 v I vT =

I - 4 v vT + 4 v vT = I

quod erat demonstrandum.